OpenStudy (angelwings996):

Need Help with D... Suppose you drive an average of 15 000 miles per year, and your car gets 24 miles per gallon. Suppose gasoline costs $3.60 a gallon. A. How much money do you spend each year on gasoline? MY ANSWER: 2250 B. You plan to trade in your car for one that gets x more miles per gallon. Write an expression to represent the new yearly cost of gasoline. MY ANSWER: 1500/24+x (3.60) C. Write an expression to represent your total savings on gasoline per year. MY ANSWER: 2250-(1500/24+x) (3.60) D. Suppose you save $600 a year with the new car. How many miles per gallon does the new

OpenStudy (angelwings996):

B. \[\frac{ 1500 }{ 24 + x }(3.6)\] C. \[2250 - \frac{ 1500 }{ 24+x }(3.6)\]

OpenStudy (aripotta):

use the expression you made for b and make it equal to 2250 - 600 1500/24+x (3.60) = 1650 so now you should be solving for x

OpenStudy (aripotta):

then you can do 24 + x to get your answer once you've figured out x

OpenStudy (angelwings996):

Oh okay, thank you!!

OpenStudy (aripotta):

no problem